# Statistical Control of Relaxation and Synchronization in Open Anyonic Systems

### v2

@@ Appendix B Anyonic Algebra for the 2-oscillator system
$[\tilde{b}_{\pm},\tilde{b}_{\pm}^{\dagger}]\approx\frac{1}{2}\left(e^{i\theta N_{1}}+e^{i\theta N_{2}}\right),$ (30)
@@
$L^{\pm}_{\text{em}}=\sqrt{\gamma(n_{\theta}+1)}\sqrt{1\pm\xi}\cdot\frac{1}{2}\left(\tilde{b}_{+}\pm e^{-i\theta/2}\tilde{b}_{-}\right),$ (36)
@@
$L^{\pm}_{\text{abs}}=\sqrt{\gamma n_{\theta}}\sqrt{1\pm\xi}\cdot\frac{1}{2}\left(\tilde{b}_{+}\pm e^{-i\theta/2}\tilde{b}_{-}\right).$ (37)
@@
$L^{\pm}_{\text{em}}$ $=\sqrt{\gamma(n_{\theta}+1)}\sqrt{1\pm\xi}\cdot\frac{1}{2}\left(\tilde{b}_{+}\pm e^{-i\theta/2}\tilde{b}_{-}\right),$ (39)
$L^{\pm}_{\text{abs}}$ $=\sqrt{\gamma n_{\theta}}\sqrt{1\pm\xi}\cdot\frac{1}{2}\left(\tilde{b}_{+}\pm e^{-i\theta/2}\tilde{b}_{-}\right).$ (40)
@@
$\mathcal{D}_{k}^{\dagger}[\tilde{b}_{+}]=-\left(|\lambda_{k}^{(+)}|^{2}\tilde{b}_{+}+\lambda_{k}^{(+)}\lambda_{k}^{(-)*}\tilde{b}_{-}\right),$ (58)

### v3

@@ Appendix B Anyonic Algebra for the 2-oscillator system
$\tilde{b}_{\pm},\tilde{b}_{\pm}^{\dagger}]\approx\frac{1}{2}\left(e^{i\theta N_{1}}+e^{i\theta N_{2}}\right),$ (32)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Appendix C Computational Notebook
The accompanying Mathematica notebook linked in the Data Availability Statement contains the results of our symbolic derivations and numerical simulations corresponding to the data presented in Figures 1–3 of the main manuscript. The file includes parameter sweeps over the statistical angle $\theta$, dissipation rate calculations, and spectral response simulations based on the third-order susceptibility. Mathematica notebooks for the figures are available online at this Wolfram Cloud link. An accompanying WolframLanguage package is available as here.